# On the Exact Linearization of Minimally Underactuated Configuration Flat Lagrangian Systems in Generalized State Representation

### v2

@@ 1 Introduction
The concept of differential flatness, as introduced, e.g., in Fliess et al. (1995), has significantly advanced nonlinear control theory. Roughly speaking, a nonlinear control system in classical state representation $\dot{x}=f(x,u)$ is flat if the $n$-dimensional classical state $x$ and the $m$-dimensional control input $u$ can be expressed in terms of an $m$-tuple $y$ of differentially independent functions $y_{j}=\varphi_{j}(x,u,u^{(1)},\ldots,u^{(\nu)})$, $j=1,\ldots,m$, and their time derivatives, where $u^{(\nu)}$ denotes the $\nu$-th time derivative of the input. However, computing a so-called flat output $y$ continues to be a complex challenge lacking a comprehensive solution, as emphasized in recent studies such as Gstöttner et al. (2023), Gstöttner et al. (2023a), Nicolau and Respondek (2019), Nicolau and Respondek (2017), and Schöberl and Schlacher (2014). Exploiting the property of flatness enables elegant and systematic trajectory planning and tracking solutions. The concept of flatness-based tracking control design typically involves exact linearization, where a nonlinear control system is exactly linearized by endogenous feedback, resulting in a closed-loop system consisting of $m$ integrator chains $y_{j}^{(\kappa_{j})}=w_{j}$, $j=1,\ldots,m$, with respective lengths $\kappa_{j}$ between a new input $w$ and the flat output $y$. As shown, e.g., in Fliess et al. (1999), the exact linearization of flat systems is systematically achievable by applying dynamic feedback laws. However, this approach has the disadvantage of increasing the order of the closed-loop system dynamics, consequently resulting in an increased order of the tracking error dynamics. As an alternative, Delaleau and Rudolph (1998) show that every flat system is linearizable by a quasi-static feedback law $u=\alpha(\tilde{x}_{B},w,w^{(1)},\ldots)$ of a so-called generalized Brunovský state $\tilde{x}_{B}$ that consists of suitably chosen consecutive time derivatives of the flat output $y$ meeting $\mathrm{dim}(\tilde{x}_{B})=\mathrm{dim}(x)$. Consequently, applying such a feedback law preserves the order of the closed-loop system dynamics. However, practically measuring or determining $\tilde{x}_{B}$ from available measurement data is often rather complex. Hence, the exact linearization of nonlinear control systems by a quasi-static feedback law $u=\alpha(x,w,w^{(1)},\ldots)$ of the classical state $x$ is of significant interest. For systems in classical state representation with a flat output of the form $y=\varphi(x,u)$, known as $(x,u)$-flat systems, Delaleau and da Silva (1998) demonstrate in a differential algebraic setting that a linearizing feedback law $u=\alpha(x,w,w^{(1)},\ldots)$ is guaranteed to exist. In a differential geometric framework, Gstöttner et al. (2023b) derive similar results and present an algorithm that constructs the lengths of the integrator chains of the closed-loop system using successive coordinate transformations, avoiding Brunovský states. Kolar et al. (2022) extend these results to differentially flat discrete-time systems.
@@
This paper discusses the exact linearization of minimally underactuated configuration flat generalized LC systems by quasi-static feedback of its generalized states. These generalized states consist of those configuration variables $\tilde{q}$ that are not selected as control inputs and their respective velocities $\tilde{v}$. Our approach builds upon essential results outlined in Gstöttner et al. (2023b) and Hartl et al. (2024), and extends them to generalized LC systems. Based on the Jacobian of the parameterization of the system variables $(\tilde{q},\tilde{v})$ by the flat output $y$, we demonstrate how to choose suitable orders of the time derivatives of the flat output intended as new inputs. We present all possible lengths of integrator chains that can be achieved by applying a quasi-static feedback law of the generalized state $(\tilde{q},\tilde{v})$ that exhibits no singularities at equilibrium points. Our approach uses neither successive coordinate transformations nor Brunovský states. The structure of this paper is organized as follows: Section 2 establishes the notation. Section 3 examines properties of differentially flat minimally underactuated classical LC systems. Section 4 covers the principles of differential flatness and exact linearization for generalized state systems. Section 5 formally introduces generalized LC systems. Finally, Section 6 is devoted to our main results, accompanied by a practical example.
@@ 4 Flatness of Generalized Systems
The objective of exact linearization is to establish a linear input-output behavior $y^{j}_{[\kappa^{j}]}=w^{j},j=1,\ldots,m,$ between a newly introduced input $w$ and the flat output $y$ by applying a suitable feedback law. However, a major challenge lies in choosing the appropriate lengths $\kappa$ of the integrator chains. In the context of tracking-control design, it is beneficial to reduce the orders of the tracking error dynamics by selecting the orders $\kappa^{j}$ as low as possible, which can be achieved by constructing a linearizing quasi-static feedback law of the generalized state $\tilde{x}$. In the following, we present conditions that ensure that selected time derivatives of a flat output can be introduced as new closed-loop input $w$ by applying a quasi-static feedback law of $\tilde{x}$. With the focus on quasi-static feedback of $\tilde{x}$, we extend Theorem 3.1 from Gstöttner et al. (2023b) to generalized state systems (11).
@@
Consider system (11) with a flat output $y=\varphi(\tilde{x},u,u_{[1]},\ldots)$. An $m$-tuple of time derivatives of this flat output given by $y_{[\kappa]}=\varphi_{[\kappa]}(\tilde{x},u,u_{[1]},\ldots)$, with $\kappa\leq R$ and #$\#\kappa=n$, can be introduced as a new input $w$ by a quasi-static feedback law $u=\alpha(\tilde{x},w,w_{[1]},\ldots)$ of the generalized state $\tilde{x}$ if and only if the differentials $\mathrm{d}\tilde{x},\mathrm{d}\varphi_{[\kappa]},\mathrm{d}\varphi_{[\kappa+1]},\ldots$ are linearly independent. [^4] The proof is omitted here for brevity but can be conducted analogously to the proof of Theorem 3.1 in Gstöttner et al. (2023b).
@@ 6.1 Exact linearization of the planar aerial manipulator
Since $\partial_{y}F_{\tilde{q}}\big{|}_{y_{s}}=I_{3\times 3}$ holds, case i) of Theorem 5 applies. Given $p-k=3$, the permutation (27) is clearly given by $y=\bar{y}=(y^{1},y^{2},y^{3})$. Hence, there exists a linearizing quasi-static feedback law (22) that is well-defined around equilibrium points $y_{s}$. The closed-loop system (23) exhibits a linear input-output behavior (24) with $\tilde{\kappa}_{1}=(2,2,2)$.
Alternatively, according to case ii), we can select $\bar{y}=(y^{2},y^{3})$ and have $\text{rank}(\partial_{\bar{y}}F_{\tilde{q}})\big{|}_{y_{s}}=p-k-1=2$. Next, we choose the column $\partial_{y^{3}_{[2]}}F_{\tilde{q}}\big{|}_{y_{s}}$ of (39) such that
@@ 7 Conclusion
Our approach illustrates a straightforward method for determining the integrator chain lengths $\tilde{\kappa}$ of an exactly linearized minimally underactuated generalized LC system with a configuration flat output. The obtained quasi-static feedback (22) of the generalized state $(\tilde{q},\tilde{v})$ then allows for rest-to-rest transitions and can be systematically derived without requiring the use of successive coordinate transformations or Brunovský states. The presented approach is suitable for any minimally underactuated configuration flat generalized LC system. However, note that neither the existence of a quasi-static feedback law (22) that is regular at equilibrium points nor the existence of a quasi-static feedback law (22), in general, is guaranteed.

### v3

@@ 1 Introduction
The concept of differential flatness, as introduced, e.g., in Fliess et al. (1995), has significantly advanced nonlinear control theory. Roughly speaking, a nonlinear control system in classical state representation $\dot{x}=f(x,u)$ is flat if the $n$-dimensional classical state $x$ and the $m$-dimensional control input $u$ can be expressed in terms of an $m$-tuple $y$ of differentially independent functions $y_{j}=\varphi_{j}(x,u,u^{(1)},\ldots,u^{(\nu)})$, $j=1,\ldots,m$, and their time derivatives, where $u^{(\nu)}$ denotes the $\nu$-th time derivative of the input. However, computing a so-called flat output $y$ continues to be a complex challenge lacking a comprehensive solution, as emphasized in recent studies such as Gstöttner et al. (2023), Gstöttner et al. (2023), Nicolau and Respondek (2019), Nicolau and Respondek (2017), and Schöberl and Schlacher (2014). Exploiting the property of flatness enables elegant and systematic trajectory planning and tracking solutions. The concept of flatness-based tracking control design typically involves exact linearization, where a nonlinear control system is exactly linearized by endogenous feedback, resulting in a closed-loop system consisting of $m$ integrator chains $y_{j}^{(\kappa_{j})}=w_{j}$, $j=1,\ldots,m$, with respective lengths $\kappa_{j}$ between a new input $w$ and the flat output $y$. As shown, e.g., in Fliess et al. (1999), the exact linearization of flat systems is systematically achievable by applying dynamic feedback laws. However, this approach has the disadvantage of increasing the order of the closed-loop system dynamics, consequently resulting in an increased order of the tracking error dynamics. As an alternative, Delaleau and Rudolph (1998) show that every flat system is linearizable by a quasi-static feedback law $u=\alpha(\tilde{x}_{B},w,w^{(1)},\ldots)$ of a so-called generalized Brunovský state $\tilde{x}_{B}$ that consists of suitably chosen consecutive time derivatives of the flat output $y$ meeting $\mathrm{dim}(\tilde{x}_{B})=\mathrm{dim}(x)$. Consequently, applying such a feedback law preserves the order of the closed-loop system dynamics. However, practically measuring or determining $\tilde{x}_{B}$ from available measurement data is often rather complex. Hence, the exact linearization of nonlinear control systems by a quasi-static feedback law $u=\alpha(x,w,w^{(1)},\ldots)$ of the classical state $x$ is of significant interest. For systems in classical state representation with a flat output of the form $y=\varphi(x,u)$, known as $(x,u)$-flat systems, Delaleau and da Silva (1998) demonstrate in a differential algebraic setting that a linearizing feedback law $u=\alpha(x,w,w^{(1)},\ldots)$ is guaranteed to exist. In a differential geometric framework, Gstöttner et al. (2024) derive similar results and present an algorithm that constructs the lengths of the integrator chains of the closed-loop system using successive coordinate transformations, avoiding Brunovský states. Kolar et al. (2022) extend these results to differentially flat discrete-time systems.
@@
This paper discusses the exact linearization of minimally underactuated configuration flat generalized LC systems by quasi-static feedback of its generalized states. These generalized states consist of those configuration variables $\tilde{q}$ that are not selected as control inputs and their respective velocities $\tilde{v}$. Our approach builds upon essential results outlined in Gstöttner et al. (2024) and Hartl et al. (2024), and extends them to generalized LC systems. Based on the Jacobian of the parameterization of the system variables $(\tilde{q},\tilde{v})$ by the flat output $y$, we demonstrate how to choose suitable orders of the time derivatives of the flat output intended as new inputs. We present all possible lengths of integrator chains that can be achieved by applying a quasi-static feedback law of the generalized state $(\tilde{q},\tilde{v})$ that exhibits no singularities at equilibrium points. Our approach uses neither successive coordinate transformations nor Brunovský states. The structure of this paper is organized as follows: Section 2 establishes the notation. Section 3 examines properties of differentially flat minimally underactuated classical LC systems. Section 4 covers the principles of differential flatness and exact linearization for generalized state systems. Section 5 formally introduces generalized LC systems. Finally, Section 6 is devoted to our main results, accompanied by a practical example.
@@ 4 Flatness of Generalized Systems
The objective of exact linearization is to establish a linear input-output behavior $y^{j}_{[\kappa^{j}]}=w^{j},j=1,\ldots,m,$ between a newly introduced input $w$ and the flat output $y$ by applying a suitable feedback law. However, a major challenge lies in choosing the appropriate lengths $\kappa$ of the integrator chains. In the context of tracking-control design, it is beneficial to reduce the orders of the tracking error dynamics by selecting the orders $\kappa^{j}$ as low as possible, which can be achieved by constructing a linearizing quasi-static feedback law of the generalized state $\tilde{x}$. In the following, we present conditions that ensure that selected time derivatives of a flat output can be introduced as new closed-loop input $w$ by applying a quasi-static feedback law of $\tilde{x}$. With the focus on quasi-static feedback of $\tilde{x}$, we extend Theorem 3.1 from Gstöttner et al. (2024) to generalized state systems (11).
@@
Consider system (11) with a flat output $y=\varphi(\tilde{x},u,u_{[1]},\ldots)$. An $m$-tuple of time derivatives of this flat output given by $y_{[\kappa]}=\varphi_{[\kappa]}(\tilde{x},u,u_{[1]},\ldots)$, with $\kappa\leq R$ and #$\#\kappa=n$, can be introduced as a new input $w$ by a quasi-static feedback law $u=\alpha(\tilde{x},w,w_{[1]},\ldots)$ of the generalized state $\tilde{x}$ if and only if the differentials $\mathrm{d}\tilde{x},\mathrm{d}\varphi_{[\kappa]},\mathrm{d}\varphi_{[\kappa+1]},\ldots$ are linearly independent. [^4] The proof is omitted here for brevity but can be conducted analogously to the proof of Theorem 3.1 in Gstöttner et al. (2024).
@@ 6.1 Exact linearization of the planar aerial manipulator
Since $\partial_{y}F_{\tilde{q}}|_{y_{s}}=I_{3\times 3}$ holds, case i) of Theorem 5 applies. Given $p-k=3$, the permutation (27) is clearly given by $y=\bar{y}=(y^{1},y^{2},y^{3})$. Hence, there exists a linearizing quasi-static feedback law (22) that is well-defined around equilibrium points. The closed-loop system (23) exhibits a linear input-output behavior (24) with $\tilde{\kappa}_{1}=(2,2,2)$.
Alternatively, according to case ii), we can select $\bar{y}=(y^{2},y^{3})$ and have $\text{rank}(\partial_{\bar{y}}F_{\tilde{q}}|_{y_{s}})=p-k-1=2$. Next, we choose the column $\partial_{y^{3}_{[2]}}F_{\tilde{q}}|_{y_{s}}$ of (39) such that
@@ 7 Conclusion
Our approach illustrates a straightforward method for determining the integrator chain lengths $\tilde{\kappa}$ of an exactly linearized minimally underactuated generalized LC system with a configuration flat output. The obtained quasi-static feedback (22) of the generalized state $(\tilde{q},\tilde{v})$ then allows for rest-to-rest transitions and can be systematically derived without requiring the use of successive coordinate transformations or Brunovský states. The presented approach is suitable for any minimally underactuated configuration flat generalized LC system. However, note that neither the existence of a quasi-static feedback law (22) that is regular around equilibrium points nor the existence of a quasi-static feedback law (22), in general, is guaranteed.